O(0,0) 4. Two small metal spheres initially neutral are placed 3.0 m apart. Now one billion electrons are transferred from one sphere to the other. Determine the charge on each sphere and the force between them.
Answer #1

here,

number of electrons , n = 1 billion = 10^9 electrons

the charge on each sphere , q = 10^9 * 1.6 * 10^-19 C

q = 1.6 * 10^-10 C

the charge on each sphere is - (1.6 * 10^-10 C) and (1.6 * 10^-10 C)

the force between them , F = k * q^2 /r^2

F = 9* 10^9 * (1.6 * 10^-10)^2 /(3^2)

F = 2.56 * 10^-11 N